a bus starting from rest moves with a uniform acceleration of 0.1 m s 2 for 2 minutes.find (a) the speed speed acquired (by) the distance travelled
1
Expert's answer
2012-07-20T11:39:32-0400
(a) the speed acquired

V = a·t = 0.1 m/s² · 120 s = 12 m/s.

(b) the distance travelled

D = a·t²/2 = 0.1 m/s² · (120 s)² / 2 = 720 m.